# Context-aware Decoding Reduces Hallucination in Query-focused Summarization

### v2

@@ Abstract
Query-focused summarization (QFS) aims to provide a summary of a single document/multi documents that can satisfy the information needs of a given query. It is useful for various real-world applications, such as abstractive snippet generation or more recent retrieval augmented generation (RAG). A prototypical QFS pipeline consists of a retriever (sparse or dense retrieval) and a generator (usually a large language model). However, applying large language models (LLM) potentially leads to hallucinations, especially when the evidence contradicts the prior belief of LLMs. There has been growing interest in developing new decoding methods to improve generation quality and reduce hallucination. In this work, we conduct a large-scale reproducibility study on one recently proposed decoding method—Context-aware Decoding (CAD). In addition to replicating CAD’s experiments on news summarization datasets, we include experiments on QFS datasets, and conduct more rigorous analysis on computational complexity and hyperparameter sensitivity. Experiments with eight different language models show that performance-wise, CAD improves QFS quality by (1) reducing factuality errors/hallucinations while (2) mostly retaining the match of lexical patterns, measured by ROUGE scores, while also at a cost of increased inference-time FLOPs and reduced decoding speed. The code implementation based on Huggingface Library is made available  [^1] https://github.com/zhichaoxu-shufe/context-aware-decoding-qfs..
@@ 1 Introduction
Query-focused summarization (QFS) aims to provide a summary of a single document/multiple documents that can satisfy the information need of a given query [XC23, Mar06]. It is useful for various real-world applications, such as abstractive snippet generation [Che+20] or more recent retrieval augmented generation (RAG) [Lew+20]. A prototypical QFS pipeline consists of a retriever to retrieve the documents to use as evidence, and a generator (usually based on large language models) to read the documents and generate natural language answers/summaries.
While the idea of combining large language models (LLM) and information retrieval seems tempting [Met+21], a few problems persist. For example, deploying generative models requires more inference FLOPs and specialized hardwares (GPUs or TPUs); and more importantly, the usage of LLMs in QFS/RAG pipeline also potentially leads to the hallucination problem where the generated summary contains information contradicting the source documents/evidence [SB22, Ben+21], which might hurt users’ experience in information seeking processes and potentially escalate the spread of misinformation. As such, mainstream search engines (Google, Bing, etc.) still deploy extractive snippets [BC14].
In parallel to training better language models to reduce hallucinations, there has been growing interest in developing decoding methods to allow the language models to better attend to evidence or to penalize degeneration, thus improving generation quality. For example, [Li+23, ] propose contrastive decoding, leveraging the output probability distribution difference between an expert LM and an amateur LM to improve generation quality. DExperts [Liu+21] focus on reducing toxicity by contrasting the expert LM with toxicity LM. [XC23, ] show that directly injecting knowledge constraints can potentially produce more faithful summarization. A few works [Shi+23, PCM22] make use of pointwise mutual information to let the LM focus more on the input context to reduce hallucination.
In this work, we focus on Context-aware Decoding [Shi+23] (CAD). CAD leverages the idea of pointwise mutual information and proposes a product-of-experts enhancement to the vanilla decoding method to make the generation more conditioned on the input evidence. We conduct a large-scale reproducibility study with different choices of language models to study CAD’s effectiveness in query-focused summarization, and include more rigorous analysis on computational complexity and hyperparameter sensitivity. Extensive experiments show that CAD improves QFS quality by (1) reducing factual mistakes/hallucinations while (2) mostly retaining the match of lexical patterns, measured by ROUGE scores. On the other hand, CAD also introduces additional inference-time FLOPs and potentially slows down decoding speed.
@@ 2 Background
Here $\tau$ is a temperature hyperparameter to control the sharpness of probability distribution over the candidate vocabulary $\mathcal{V}$. Examples of the sampling strategy include greedy and Top-$p$ sampling [Hol+19].
As we are operating on a conditional generation problem, we hope the generation is more conditioned on input context $\mathbf{c}$. [Shi+23] utilize the idea of pointwise mutual information (PMI). We consider two events $\mathbf{c}$ and $y_{t}$:
@@
The interpretation of the above equation is we would like to measure the association of event $y_{t}$ (predicting a specific token) and event $\mathbf{c}$ (the presence of context $\mathbf{c}$). $p_{\theta}(y_{t}|\mathbf{x},\mathbf{y}_{<t})$ is the prior probability, representing the model’s prior belief from its parameters $\theta$ when no presence of context $\mathbf{c}$, whereas $p_{\theta}(y_{t}|\mathbf{c},\mathbf{x},\mathbf{y}_{<t})$ is the posterior probability with $\mathbf{c}$. To leverage PMI, [Shi+23, ] propose to multiply PMI with the original probability distribution:
@@
Here larger $\alpha$ yields more weight on the context $\mathbf{c}$ and $\alpha=0$ reduces to vanilla decoding. [Shi+23] recommend setting $\alpha=0.5$ for summarization tasks and we also examine this choice in Section 3.
@@
We follow the notations used in [Kap+20]. Let us consider the Transformer architecture with following hyperparameters: number of layers $n_{\text{layer}}$, dimension of the embedding layer and intermediate representation $d_{\text{model}}$, dimension of the intermediate feedforward layer $d_{\text{ff}}$, dimension of the attention output $d_{\text{attn}}$ and number of attention heads $n_{\text{heads}}$; and denote the number of tokens in the input sequence as $n_{\text{input}}$. The model size $N$ is the number of non-embedding parameters of the model, $N\approx|\theta|$:
@@
The total FLOPs per token in one forward pass is approximately $C_{\text{forward}}=2N+2n_{\text{layer}}\,n_{\text{input}}\,d_{\text{attn}}$ (see Section 2.1 in [Kap+20] for details). In LLMs, $d_{\text{model}}\gg n_{\text{input}}/12$, we can roughly ignore the second term and use $C_{\text{forward}}\approx 2N$.
@@ 3.2 Experiment Setup
[…]
@@
  • LLaMA-7B [Tou+23] is a decoder-only language model pre-trained with causal language modeling objective from FAIR.
  • Llama-2-7B [Tou+23] is the second iteration of LLaMA model by better pre-training corpus and minor architectural improvements.
  • MPT-7B [Tea23] is a decoder-only language model pre-trained with causal language modeling objective from MosaicML
@@
  • Mistral-7B-v0.1 [Jia+23] is a decoder-only language model pre-trained with causal language modeling objective from Mistral AI.
@@
  • Flan-T5-xl [Chu+22] (3B) is an encoder-decoder model from T5 family [Raf+20]; it is pre-trained with span-infilling objective and further adapted with causal language modeling objective. It is instruction finetuned on the Flan collection.
  • Flan-T5-xxl [Chu+22] (11B) undergoes same training procedure as Flan-T5-xl and demonstrates strong performance on natural language understanding tasks.
@@
Similar to [Shi+23], we evaluate the generation quality with ROUGE F1 [Lin04], BERTScore-precision [Zha+19] and FactKB [Fen+23]. ROUGE measures the match of lexical patterns while BERTScore measures the semantic similarity between the generation and the reference. FactKB is a model-based evaluation metric that classifies whether the generation is faithful to the document. For BERTScore we use roberta-large-mnli as it shows a higher correlation with human judgments. For FactKB we use the original classifier trained by the authors [Fen+23].
@@
Our experimental settings follow [Shi+23] to a large extent. For decoding, we set temperature=1, use a combination of Top-$k$ and Top-$p$ sampling, and vary $\alpha$ to study the effectiveness of CAD. For the main results reported in Section 4, we use $\alpha=0.3$ and use 0.15, 0.5 as an ablation study. Detailed choice of hyperparameters can be found in Appendix 7.2.
@@ 4 Results and Analysis
  • RQ1: Performance-wise, can CAD (1) improve news summarization performance as discussed by [Shi+23]? and (2) reduce hallucination/factuality errors in QFS?
@@
We present the results on QFS datasets and news summarization datasets in Table 2 and Table 1, respectively. On news summarization datasets, CAD achieves improved ROUGE scores with all eight models on CNN Dailymail dataset, and improves with six models on XSUM datasets. In terms of factuality errors, CAD improves FactKB scores on six out of eight models on CNN Dailymail dataset and eight out of eight models on XSUM dataset. These results support the finding in [Shi+23] that CAD can improve news summarization performance.
@@
Datasets $\bm{\alpha}$ value ROUGE-1 ROUGE-2 ROUGE-L BERTScore-P FactKB
@@ 5 Related Work
[Mar06] categorize search into three two different information seeking processes—Information Lookup and Exploratory Search. Information lookup refers to the user looking for simple facts or asking factoid questions, while exploratory search/search to learn often involves multiple iterations. Different approaches have been explored to reduce users’ cognitive loads in the information-seeking process. For example, Perplexity AI [^3] https://www.perplexity.ai/ transfers web search queries into QA-style questions, and cites sources in the answers to improve verifiability. Snippets [BC14] provide short previews for each webpage to allow users to locate relevant information without examining the webpage. Counterfactual explanations [Xu+23] suggest counterfactual queries as explanations and demonstrate search sessions’ effectiveness can be improved.
@@
[Ji+23, ] present a survey of hallucination in natural language generation. There are mainly two types of hallucinations, namely intrinsic hallucination and extrinsic hallucination. Intrinsic hallucination refers to the generated output that contradicts the source content, whereas extrinsic hallucination refers to the generated output that cannot be verified from the source content. In retrieval augmented generation, we mainly focus on intrinsic hallucination. Pre-trained language models (not instruction finetuned) are more prone to hallucination, as the gigantic pre-training corpus may contain untruthful texts and the teacher forcing training procedure lets the models pick up this untruthful information. [Shu+21, ] show that retrieval augmentation can reduce hallucinations in downstream tasks. [Tia+23, ] show that language models can be trained to reduce hallucination via reinforcement learning techniques. [Var+23, ] propose to use lm logits to detect hallucination in the generation process and mitigate it by retrieving evidence and rewriting.
@@
[Li+23, ] propose Contrastive Decoding (CD) to improve generation quality. The idea is to contrast the stronger expert model’s prediction with a smaller amateur model. A recent work [OL23] shows that CD can also improve the reasoning ability of the language models on multiple reasoning benchmark datasets. Based on a similar idea, [Liu+21, ] show that by contrasting to a toxic model (language model trained with toxicity contents), the toxicity of the expert language model’s generations can be reduced. A different flavor of controlled generation aims to let the model generate specific tokens/phrases. [Lu+21, ] show that such a goal can be achieved by utilizing specific data structure and interpolation of lm logits.

### v3

@@ Abstract
Query-focused summarization (QFS) aims to provide a summary of a single document/multi documents that can satisfy the information needs of a given query. It is useful for various real-world applications, such as abstractive snippet generation or more recent retrieval augmented generation (RAG). A prototypical QFS pipeline consists of a retriever (sparse or dense retrieval) and a generator (usually a large language model). However, applying large language models (LLM) potentially leads to hallucinations, especially when the evidence contradicts the prior belief of LLMs. There has been growing interest in developing new decoding methods to improve generation quality and reduce hallucination. In this work, we conduct a large-scale reproducibility study on one recently proposed decoding method — Context-aware Decoding (CAD). In addition to replicating CAD’s experiments on news summarization datasets, we include experiments on QFS datasets, and conduct more rigorous analysis on computational complexity and hyperparameter sensitivity. Experiments with eight different language models show that performance-wise, CAD improves QFS quality by (1) reducing factuality errors/hallucinations while (2) mostly retaining the match of lexical patterns, measured by ROUGE scores, while also at a cost of increased inference-time FLOPs and reduced decoding speed. The code implementation based on Huggingface Library is made available
@@ 1 Introduction
Query-focused summarization (QFS) aims to provide a summary of a single document/multiple documents that can satisfy the information need of a given query (Xu and Cohen, 2023; Marchionini, 2006). It is useful for various real-world applications, such as abstractive snippet generation (Chen et al., 2020) or more recent retrieval augmented generation (RAG) (Lewis et al., 2020; Gao et al., 2024). A prototypical QFS pipeline consists of a retriever to retrieve the documents to use as evidence, and a generator (usually based on large language models) to read the documents and generate natural language answers/summaries.
While the idea of combining large language models (LLM) and information retrieval seems tempting (Metzler et al., 2021), a few problems persist. For example, deploying generative models requires more inference FLOPs and specialized hardwares (GPUs or TPUs); and more importantly, the usage of LLMs in QFS/RAG pipeline also potentially leads to the hallucination problem where the generated summary contains information contradicting the source documents/evidence (Shah and Bender, 2022; Bender et al., 2021), which might hurt users’ experience in information seeking processes and potentially escalate the spread of misinformation. As such, mainstream search engines (Google, Bing, etc.) still deploy extractive snippets (Bast and Celikik, 2014).
In parallel to training better language models to reduce hallucinations, there has been growing interest in developing decoding methods to allow the language models to better attend to evidence or to penalize degeneration, thus improving generation quality. For example, Li et al. (2023) propose contrastive decoding, leveraging the output probability distribution difference between an expert LM and an amateur LM to improve generation quality. DExperts (Liu et al., 2021) focus on reducing toxicity by contrasting the expert LM with toxicity LM. Xu and Cohen (2023) show that directly injecting knowledge constraints can potentially produce more faithful summarization. A few works (Shi et al., 2023; van der Poel et al., 2022) make use of pointwise mutual information to let the LM focus more on the input context to reduce hallucination.
In this work, we focus on Context-aware Decoding (CAD, Shi et al., 2023). CAD leverages the idea of pointwise mutual information and proposes a product-of-experts enhancement to the vanilla decoding method to make the generation more conditioned on the input evidence. We conduct a large-scale reproducibility study with different choices of language models to study CAD’s effectiveness in query-focused summarization, and include more rigorous analysis on computational complexity and hyperparameter sensitivity. Extensive experiments show that CAD improves QFS quality by (1) reducing factual mistakes/hallucinations while (2) mostly retaining the match of lexical patterns, measured by ROUGE scores. On the other hand, CAD also introduces additional inference-time FLOPs and potentially slows down decoding speed.
@@ 2 Background
Here $\tau$ is a temperature hyperparameter to control the sharpness of the probability distribution over the candidate vocabulary $\mathcal{V}$. Examples of the sampling strategy include greedy and Top-$p$ sampling (Holtzman et al., 2019).
As we are operating on a conditional generation problem, we hope the generation is more conditioned on input context $\mathbf{c}$. Shi et al. (2023) utilize the idea of pointwise mutual information (PMI). We consider two events $\mathbf{c}$ and $y_{t}$:
@@
The interpretation of the above equation is we would like to measure the association of event $y_{t}$ (predicting a specific token) and event $\mathbf{c}$ (the presence of context $\mathbf{c}$). $p_{\theta}(y_{t}|\mathbf{x},\mathbf{y}_{<t})$ is the prior probability, representing the model’s prior belief from its parameters $\theta$ when no presence of context $\mathbf{c}$, whereas $p_{\theta}(y_{t}|\mathbf{c},\mathbf{x},\mathbf{y}_{<t})$ is the posterior probability with $\mathbf{c}$. To leverage PMI, Shi et al. (2023) propose to multiply PMI with the original probability distribution:
@@
Here larger $\alpha$ yields more weight on the context $\mathbf{c}$ and $\alpha=0$ reduces to vanilla decoding. Shi et al. (2023) recommend setting $\alpha=0.5$ for summarization tasks and we also examine this choice in Section 3.
@@
We follow the notations used in Kaplan et al. (2020). Let us consider the Transformer architecture with following hyperparameters: number of layers $n_{\text{layer}}$, dimension of the embedding layer and intermediate representation $d_{\text{model}}$, dimension of the intermediate feedforward layer $d_{\text{ff}}$, dimension of the attention output $d_{\text{attn}}$ and number of attention heads $n_{\text{heads}}$; and denote the number of tokens in the input sequence as $n_{\text{input}}$. The model size $N$ is the number of non-embedding parameters of the model, $N\approx|\theta|$:
@@
The total FLOPs per token in one forward pass is approximately $C_{\text{forward}}=2N+2n_{\text{layer}}\,n_{\text{input}}\,d_{\text{attn}}$ (see Section 2.1 in Kaplan et al. (2020) for details). In LLMs, $d_{\text{model}}\gg n_{\text{input}}/12$, we can roughly ignore the second term and use $C_{\text{forward}}\approx 2N$.
@@ 3.2 Experiment Setup
We benchmark CAD’s performance on two widely adopted QFS datasets. Debatepedia dataset is collected by Nema et al. (2017) from 663 debates of 53 diverse categories in an encyclopedia of debates and consists of 12K/0.7K/1.0K query-document summarization triplets $(\emph{query},\emph{document},\emph{summarization})$. PubMedQA (Jin et al., 2019) is a long-form abstractive question-answering dataset from the biomedical domain with the contexts available. We also use two news summarization datasets from the original CAD paper: CNN Dailymail (See et al., 2017) and XSUM (Narayan et al., 2018) for reproducibility purposes.
@@
  • LLaMA-7B (Touvron et al., 2023a) is a decoder-only language model pre-trained with causal language modeling objective from FAIR.
  • Llama-2-7B (Touvron et al., 2023b) is the second iteration of LLaMA model by better pre-training corpus and minor architectural improvements.
  • MPT-7B (Team, 2023) is a decoder-only language model pre-trained with causal language modeling objective from MosaicML
@@
  • Mistral-7B-v0.1 (Jiang et al., 2023) is a decoder-only language model pre-trained with causal language modeling objective from Mistral AI.
@@
  • Flan-T5-xl (Chung et al., 2022) (3B) is an encoder-decoder model from T5 family Raffel et al. (2020); it is pre-trained with span-infilling objective and further adapted with the causal language modeling objective. It is instruction finetuned on the Flan collection.
  • Flan-T5-xxl (Chung et al., 2022) (11B) undergoes same training procedure as Flan-T5-xl and demonstrates strong performance on natural language understanding tasks.
@@
Similar to Shi et al. (2023), we evaluate the generation quality with ROUGE F1 (Lin, 2004), BERTScore-precision (Zhang et al., 2019) and FactKB (Feng et al., 2023). ROUGE measures the match of lexical patterns while BERTScore measures the semantic similarity between the generation and the reference. FactKB is a model-based evaluation metric that classifies whether the generation is faithful to the document. For BERTScore we use roberta-large-mnli as it shows a higher correlation with human judgments. For FactKB we use the original classifier trained by the authors (Feng et al., 2023).
@@
Our experimental settings follow Shi et al. (2023) to a large extent. For decoding, we set temperature=1, use a combination of Top-$k$ and Top-$p$ sampling, and vary $\alpha$ to study the effectiveness of CAD. For the main results reported in Section 4, we use $\alpha=0.3$ and use 0.15, 0.5 as an ablation study. Detailed choice of hyperparameters can be found in Appendix 7.2.
@@ 4 Results and Analysis
  • RQ1: Performance-wise, can CAD (1) improve news summarization performance as discussed by Shi et al. (2023)? and (2) reduce hallucination/factuality errors in QFS?
@@
We present the results on QFS datasets and news summarization datasets in Table 2 and Table 1, respectively. On news summarization datasets, CAD achieves improved ROUGE scores with all eight models on CNN Dailymail dataset, and improves with six models on XSUM datasets. In terms of factuality errors, CAD improves FactKB scores on six out of eight models on CNN Dailymail dataset and eight out of eight models on XSUM dataset. These results support the finding in Shi et al. (2023) that CAD can improve news summarization performance.
@@
Datasets $\boldsymbol{\alpha}$ value ROUGE-1 ROUGE-2 ROUGE-L BERTScore-P FactKB
@@ 5 Related Work
Marchionini (2006) categorize search into three two different information seeking processes — Information Lookup and Exploratory Search. Information lookup refers to the user looking for simple facts or asking factoid questions, while exploratory search/search to learn often involves multiple iterations. Different approaches have been explored to reduce users’ cognitive loads in the information-seeking process. For example, Perplexity AI [^2] https://www.perplexity.ai/ transfers web search queries into QA-style questions, and cites sources in the answers to improve verifiability. Snippets (Bast and Celikik, 2014) provide short previews for each webpage to allow users to locate relevant information without examining the webpage. Counterfactual explanations (Xu et al., 2024) suggest counterfactual queries as explanations and demonstrate search sessions’ effectiveness can be improved.
@@
Ji et al. (2023) present a survey of hallucination in natural language generation. There are mainly two types of hallucinations, namely intrinsic hallucination and extrinsic hallucination. Intrinsic hallucination refers to the generated output that contradicts the source content, whereas extrinsic hallucination refers to the generated output that cannot be verified from the source content. In retrieval augmented generation, we mainly focus on intrinsic hallucination. Pre-trained language models (not instruction finetuned) are more prone to hallucination, as the gigantic pre-training corpus may contain untruthful texts and the teacher forcing training procedure lets the models pick up this untruthful information. Shuster et al. (2021) show that retrieval augmentation can reduce hallucinations in downstream tasks. Tian et al. (2023) show that language models can be trained to reduce hallucination via reinforcement learning techniques. Varshney et al. (2023) propose to use lm logits to detect hallucination in the generation process and mitigate it by retrieving evidence and rewriting.
@@
Li et al. (2023) propose Contrastive Decoding (CD) to improve generation quality. The idea is to contrast the stronger expert model’s prediction with a smaller amateur model. A recent work (O’Brien and Lewis, 2023) shows that CD can also improve the reasoning ability of the language models on multiple reasoning benchmark datasets. Based on a similar idea, Liu et al. (2021) show that by contrasting to a toxic model (language model trained with toxicity contents), the toxicity of the expert language model’s generations can be reduced. A different flavor of controlled generation aims to let the model generate specific tokens/phrases. Lu et al. (2021) show that such a goal can be achieved by utilizing specific data structure and interpolation of lm logits.